OpenStudy (anonymous):

4x=y-2 4x-y=9 solve the system by the method of your choice.

OpenStudy (anonymous):

start with \[4x-y=-2\] \[4x-y=9\] and then note that you have the same thing on the left but different numbers on the right, so there is no solution

OpenStudy (anonymous):

if x=0 then y=2; and if y=0 then x=-1/2. Keep plugging in numbers for either your x or y variables to find different solutions; assuming these are for points along a line (i.e. basic algebra)

OpenStudy (lgbasallote):

can i have a follow-up question..what if it's 0x + 0y = 0? what is that? no solution too? indeterminate? or something? @satellite73

OpenStudy (kinggeorge):

@lgbasallote I would call that indeterminate. There are infinitely many solutions, so you can't tell which one to use.

OpenStudy (lgbasallote):

oh thank you :D always got me confused

OpenStudy (anonymous):

if you get \[0x+0y=0\] then it is true for all x and y

OpenStudy (anonymous):

i.e. it is the same line